Jane travels to her grandmothers house which is 270km by car. She travels at an average speed of 50km/ hour on good roads and an average speed of 20km/hour on roads under construction. If the total trip took 6 hours, find the time it took to travel on the roads under construction, using the substitution or elimination.
Thanks

Answer by richwmiller(17219)

50*g+20*c=270
g+c=6
solve for c
g=6-c
50*g+20*c=270
50*(6-c)+20*c=270